OpenStudy (anonymous):

How do you write three equivalent fractions of 4/9

OpenStudy (jdoe0001):

just multiply both top and bottom by a same number

OpenStudy (jdoe0001):

you need 3 more, so do it 3 times

OpenStudy (anonymous):

when ur done with this guys question could you meet me back at my question please? thanks so much :D

OpenStudy (doc.brown):

\[\frac{4}{9}\times\frac{2}{2}=\frac{8}{18}\]\[\frac{4}{9}\times\frac{n}{n}=\frac{4n}{9n}\]\[\frac{4}{9}\times\frac{mile}{1.6km}=\frac{4miles}{14.4km}\]